OpenStudy (anonymous):

A triangle has one angle that measures 68° and one angle that measures 22°. What kind of triangle is it?

OpenStudy (anonymous):

The sum of all the angles has to equal 180 degrees In this case we are given two of the angles so we know: 68+22+angle3 = 180 We can find the third angle by subtracting the other two angles from 180. 180-(68+22)=angle3 can you tell me what angle 3 is?

OpenStudy (coconutjj):

right triangle 180 - 68 - 22 = 90

OpenStudy (anonymous):

thank you